Calculate the probability that the speed of a chlorine molecule will lie between 200 m s−1 and 220 m s1 at 300 K. Given that mass of chlorine molecule = 70u, kB = 1.38 × 1023 JK1 and NA = 6.02 × 1026 kmol1


1
Expert's answer
2021-04-09T11:24:41-0400

To be given in question

"v_{1}=200meter\/sec"

"v_{2}=220meter\/sec"

"M_{cl}=70 amu" "K_{B}=1.38\times10^{-23}Jule/mol" "T=300K"

To be asked in question


Probability speed

We know that

"P=[\frac{m}{2\pi K_{B}T}]^\frac{1}{2}e^{-\frac{mv^2}{2K_{B}T}}"


"P_{1}=[\frac{m}{2\pi K_{B}T}]^\frac{1}{2}e^{-\frac{mv_{1}^2}{2K_{B}T}}"

Put value

"P_{m}=[\frac{m}{2\pi K_{B}T}]^\frac{1}{2}"

"P_{m}= [\frac{70 \times 1.67\times10^{-27}}{2\times3.14\times1.38\times10^{-23}\times 300}]^\frac{1}{2}"

"P_{m}=2.12\times10^{-3}"

"\frac{mv_{1}^2}{2K_{B}T}=x"

"\frac{mv_{1}^2}{2K_{B}T}=" "[\frac{70 \times 1.67\times10^{-27}\times200^2}{2\times1.38\times10^{-23}\times 300}]^\frac{1}{2}"


"x=0.564"

"\frac{mv_{2}^2}{2K_{B}T}=y"

"\frac{mv_{2}^2}{2K_{B}T}=" "y= [\frac{70 \times 1.67\times10^{-27}\times220^2}{2\times1.38\times10^{-23}\times300}]"

"y=0.6824"

First velocity

"P_{1} =P_{m}e^{-x}"

Put value

"P_{1}=1.206\times10^{-03}"

Second velocity

"P_{2} =P_{m}e^{-y}"

Put value

"P_{2}=" "1.071\times10^{-03}"